A con is flipped 12 times what is the probability of getting
(1)no heads
2) no more than 3 heads

Answer by stanbon(75887)

A coin is flipped 12 times
Binomial Problem with n = 12 and p = 1/2
----------------------------------------
what is the probability of getting
(1)no heads
P(x = 0) = 12C0*(1/2)^0*(1/2)^12 = 1*1*(1/2)^12 = 1/2^12
---------------------------------
2) no more than 3 heads
P(0<= x <=3) = binomcdf(12,1/2,3) = 0.0730
--------------------
Cheers,
Stan H.
------------